LearnCo Sales Budget For the Year Ending December 31, 20Y2 Product Unit Sales Volume Unit Selling...

LearnCo
Sales Budget
For the Year Ending December 31, 20Y2
Product Unit Sales Volume Unit Selling Price Total Sales
Basic Abacus $8.00 $288,000
Deluxe Abacus 432,000
Totals 72,000 $720,000

X

Production Budget

The production budget should be integrated with the sales budget to ensure that production and sales are kept in balance during the year. The production budget estimates the number of units to be manufactured to meet budgeted sales and desired inventory levels.

You note that LearnCo has omitted six numbers from the following production budget and fill in the missing amounts. You may need to use numbers from the sales budget you prepared.

LearnCo
Production Budget
For the Year Ending December 31, 20Y2
Units Basic Units Deluxe
Expected units to be sold (from Sales Budget)
Desired ending inventory, December 31, 20Y2 1,000 3,000
Total units available
Estimated beginning inventory, January 1, 20Y2 (1,050) (2,100)
Total units to be produced
Answer #1
  • Requirement asked

LearnCo

Production Budget

For the Year Ending December 31, 20Y2

Units Basic

Units Deluxe

Expected units to be sold (from Sales Budget)

               36,000

         36,000

Desired ending inventory, December 31, 20Y2

                 1,000

           3,000

Total units available

               37,000

         39,000

Estimated beginning inventory, January 1, 20Y2

                (1,050)

          (2,100)

Total units to be produced

               35,950

         36,900

  • Working

LearnCo

Production Budget

For the Year Ending December 31, 20Y2

Units Basic

Units Deluxe

Expected units to be sold (from Sales Budget)

36000 [see working below]

36000 [see working below]

Desired ending inventory, December 31, 20Y2

1000

3000

Total units available

=36000+1000

=36000+3000

Estimated beginning inventory, January 1, 20Y2

-1050

-2100

Total units to be produced

=37000-1050

=39000-2100

LearnCo

Sales Budget

For the Year Ending December 31, 20Y2

Product

Unit Sales Volume

Unit Selling Price

Total Sales

Basic Abacus

               36,000 [288000 / 8]

$8

$288,000

Deluxe Abacus

               36,000 [72000 – 36000]

$12 [432000 / 36000]

$432,000

Totals

               72,000

$720,000